∫5x^4/x^5+1

The answer is In(x^5+1)+c. I know how to integrate but I don't know how to get to this answer. Can you show me the working to get to this answer?
Many Thanks
Andrew

Answer by richard1234(7193)

Just use a simple substitution and that is all. By letting u = x^5 + 1 and, du/dx = 5x^4, du = 5x^4 dx,



This is equal to

where C is a constant.